BlueCorp. is growing quickly. Dividends are expected to grow at a rate of 18 percent for the next three years, with the growth rate falling off to a constant 5.7 percent thereafter. If the required return is 9.82 percent and the company just paid a $3.36 dividend, what is the current share price?
Answer #1

HI

Here Current dividend D0 = $3.36

growth rate for next 3 years = 18%

so Dividend next year D1 = 3.36*(1+18%) = $3.97

D2 = 3.97*(1+18%) = $4.68

D3 = 4.68*(1+18%) = $5.52

then perpetual growth rate g = 5.7%

required return k = 9.82%

as per dividend discount model, a share price is sum of present values of all future dividends

So share price = 3.97/(1+9.82%) + 4.68/(1+9.82%)^2+5.52/(1+9.82%)^3+ 5.52*(1+5.7%)/(9.82%-5.7%)(1+9.82%)^3

=$118.59

Thanks